How many unordered 2-letter subsets of the letters A,B, and C are there?
Answer by stanbon(75887)

How many unordered 2-letter subsets of the letters A,B, and C are there?
----
Ordered subsets: ab ; ac ; bc-------3
Total # of 2 letter subsets: 3*2 = 6
-----
Ans: # of unordered 2-letter subsets = 6-3 = 3
========================================
Cheers,
Stan H.
============